OpenStudy (anonymous):

2sin^2 x + cos^2 x = 25/16. Find sin x.

OpenStudy (potatoes.ramu):

Try expressing sin in terms of cos or vice versa

OpenStudy (cwrw238):

use the identity cos^2 x = 1 - sin^2 x

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

@cwrw238 3/4! :)

OpenStudy (anonymous):

Thank you @cwrw238 :)

OpenStudy (cwrw238):

sin^2 x = 25/16 - 1 = 9/16 sin x = 3/4 or -3/4

OpenStudy (cwrw238):

yw